# Is the Faber-Krahn inequality true for the Stokes operator?

### v1

@@ 3.3 The ball satisfies strong second-order conditions for Problem (5)
    $\left(j_{1,1}J_{2}^{\prime}(j_{1,1})+2J_{2}(j_{1,1})\right)b_{2}+2d_{2}=-\frac{j_{1,1}}{\sqrt{\pi}}\beta_{2},$
  • in front of the term $\sin\theta$:
@@
    $-\left(j_{1,1}J_{2}^{\prime}(j_{1,1})+2J_{2}(j_{1,1})\right)a_{2}-2c_{2}=-\frac{j_{1,1}}{\sqrt{\pi}}(2\alpha_{0}-\alpha_{2}),$
  • in front of the term $\cos((n+1)\theta)$:
    $\left(nJ_{n}(j_{1,1})-j_{1,1}J_{n}^{\prime}(j_{1,1}))\right)b_{n}+\left((n+2)J_{n+2}(j_{1,1})+j_{1,1}J_{n+2}^{\prime}(j_{1,1})\right)b_{n+2}+(n+2)d_{n+2}$
    $=$ $-\frac{j_{1,1}}{\sqrt{\pi}}(\beta_{n+2}-\beta_{n}),\phantom{aaa\hskip 199.16928ptaaaaaaaaaaaaaa}$
  • in front of the term $\sin((n+1)\theta)$:
    $\left(j_{1,1}J_{n}^{\prime}(j_{1,1}-nJ_{n}(j_{1,1})))\right)a_{n}-\left((n+2)J_{n+2}(j_{1,1})+j_{1,1}J_{n+2}^{\prime}(j_{1,1})\right)a_{n+2}-(n+2)c_{n+2}$
    $=$ $-\frac{j_{1,1}}{\sqrt{\pi}}(-\alpha_{n+2}+\alpha_{n}).\phantom{aaa\hskip 199.16928ptaaaaaaaaaaaaaa}$
We proceed similarly with the function $u_{2}^{\prime}$. As $u_{2}^{\prime}=-\partial\psi_{\Phi}^{\prime}/\partial x_{1}$, it follows that
$u_{2}^{\prime}=\sum_{n=0}^{+\infty}-\cos\theta j_{1,1}J_{n}^{\prime}(j_{1,1})(a_{n}\cos(n\theta)+b_{n}\sin(n\theta))-n\sin\theta(c_{n}\cos(n\theta)+d_{n}\sin(n\theta))$
$\displaystyle+\sum_{n=0}^{+\infty}n\sin\theta J_{n}(j_{1,1})(b_{n}\cos(n\theta)-a_{n}\sin(n\theta))+n\sin\theta(d_{n}\cos(n\theta)-c_{n}\sin(n\theta))\quad\text{on }\partial{\mathbb{B}}_{2}.$
Identifying the cosine and sine terms in the relation above yields
  • in front of the constant term:
    $-j_{1,1}J_{1}^{\prime}(j_{1,1})a_{1}-c_{1}=\frac{j_{1,1}}{\sqrt{\pi}}\alpha_{1},$
  • in front of the term $\cos\theta$:
    $-\left(j_{1,1}J_{2}^{\prime}(j_{1,1})+2J_{2}(j_{1,1})\right)a_{2}-4c_{2}=\frac{j_{1,1}}{\sqrt{\pi}}(2\alpha_{0}+\alpha_{2}),$
  • in front of the term $\sin\theta$:
    $-\left(j_{1,1}J_{2}^{\prime}(j_{1,1})+2J_{2}(j_{1,1})\right)b_{2}-4d_{2}=\frac{j_{1,1}}{\sqrt{\pi}}\beta_{2},$
  • in front of the term $\cos((n+1)\theta)$:
    $\left(nJ_{n}(j_{1,1})-j_{1,1}J_{n}^{\prime}(j_{1,1}))\right)a_{n}-\left((n+2)J_{n+2}(j_{1,1})+j_{1,1}J_{n+2}^{\prime}(j_{1,1})\right)a_{n+2}-2(n+2)c_{n+2}$
    $=$ $\frac{j_{1,1}}{\sqrt{\pi}}(\alpha_{n+2}+\alpha_{n}),\phantom{aaa\hskip 199.16928ptaaaaaaaaaaaaaa}$
  • in front of the term $\sin((n+1)\theta)$:
    $\left(nJ_{n}(j_{1,1})-j_{1,1}J_{n}^{\prime}(j_{1,1}))\right)b_{n}-\left((n+2)J_{n+2}(j_{1,1})+j_{1,1}J_{n+2}^{\prime}(j_{1,1})\right)b_{n+2}-2(n+2)d_{n+2}$
    $=$ $\frac{j_{1,1}}{\sqrt{\pi}}(\beta_{n+2}+\beta_{n}).\phantom{aaa\hskip 199.16928ptaaaaaaaaaaaaaa}$
We obtain the desired expression by combining together all the above relations. ∎
@@
Recall that we assume $\varphi\in\langle 1\rangle^{\perp}$ where $1$ is the constant function and $\langle 1\rangle$ is its $L^{2}$ orthogonal subspace.
@@
(we recall that we have supposed $\varphi$ orthogonal to $1,\cos\theta,\sin\theta$, so $a_{0}=a_{1}=b_{1}=0$. Since the first positive zero of $J_{n}$ is greater than $j_{1,1}$ for all $n\geqslant 2$, one has $\gamma_{n}>0$ for every $n\geqslant 2$. Furthermore, using that $J_{n}(x)\sim\frac{1}{n!}\left(\frac{x}{2}\right)^{n}$ as $n\to+\infty$, we infer that
@@ 4 Necessary optimality conditions: proof of Theorem 4
As alluded to in the introduction, Theorem 3 follows from Theorem 4, so that it is clearer to begin with the proof of the latter (nevertheless, we give a proof of 3 in section 5 that only relies on the computations of the semi-differential analysed in the present section). This theorem relies on several results on the semi-differentiability of eigenvalues. To write things down in a precise way, we consider a domain ${\Omega}$ whose boundary is of class $\mathscr{C}^{3}$. Although $\lambda_{1}({\Omega})$ may not be simple, it has finite multiplicity, say $N$, and the map $\mathcal{F}$ is semi-shape differentiable at ${\Omega}$ [20, Theorem 2.1, Chapter 1] (see also [13, Theorem 2.1 and Remark 2.2] and [15]), in the sense that, for any vector field $\Phi\in\mathscr{X}$, the limit
$\langle\partial\mathcal{F}({\Omega}),\Phi\rangle:=\underset{t\to 0^{+}}{\lim}\frac{\mathcal{F}((\mathrm{Id}+t\Phi){\Omega})-\mathcal{F}({\Omega})}{t}$
@@
$\langle\partial\lambda_{1}({\Omega}),\Phi\rangle=-\min_{{\mathbf{u}}\in E_{1}\,,\int_{\Omega}\|{\mathbf{u}}\|^{2}=1}\int_{\partial{\mathbb{B}}_{3}}\|({\nabla}{\mathbf{u}})\nu\|^{2}\langle\Phi,\nu\rangle.$
@@
$\langle\partial\mathcal{F}({\Omega}),\Phi\rangle$ $=\frac{2}{3}\lambda_{1}({\Omega})|{\Omega}|^{\frac{1}{3}}\int_{\partial{\Omega}}\langle\Phi,\nu\rangle-|{\Omega}|^{\frac{2}{3}}\min_{{\mathbf{u}}\in E_{1}\,,\int_{\Omega}\|{\mathbf{u}}\|^{2}=1}\int_{\partial{\Omega}}\|({\nabla}{\mathbf{u}})\nu\|^{2}\langle\Phi,\nu\rangle$
$=\min_{{\mathbf{u}}\in E_{1}\,,\int_{\Omega}\|{\mathbf{u}}\|^{2}=1}\int_{\partial{\Omega}}\left(\frac{2}{3}\lambda_{1}({\Omega})|{\Omega}|^{\frac{1}{3}}-|{\Omega}|^{\frac{2}{3}}\cdot\|({\nabla}{\mathbf{u}})\nu\|^{2}\right)\langle\Phi,\nu\rangle$
$=\min_{(\alpha_{i})_{i=1,\dots,N},\ \sum_{i=1}^{N}{\alpha_{i}^{2}}=1}\sum_{i,j=1}^{N}\int_{\partial{\Omega}}\left(\frac{2}{3}\lambda_{1}({\Omega})|{\Omega}|^{\frac{1}{3}}\alpha_{i}^{2}-|{\Omega}|^{\frac{2}{3}}\alpha_{i}\alpha_{j}\langle({\nabla}{\mathbf{u}}_{i})\nu,({\nabla}{\mathbf{u}}_{j})\nu\rangle\right)\langle\Phi,\nu\rangle.$
@@
$M_{\Phi}:=\frac{2}{3}\lambda_{1}({\Omega})|{\Omega}|^{\frac{1}{3}}\int_{\partial{\Omega}}\langle\Phi,\nu\rangle\mathrm{I}_{3}-|{\Omega}|^{\frac{2}{3}}\left(\int_{\partial{\Omega}}\langle({\nabla}{\mathbf{u}}_{i})\nu,({\nabla}{\mathbf{u}}_{j})\nu\rangle\langle\Phi,\nu\rangle\right)_{1\leqslant i,j\leqslant N}.$
@@
We thus obtain the following fact: if ${\Omega}^{*}$ is optimal then for any $(i,j)\in\{1,\dots,N\}^{2}$, for any $x\in\partial{\Omega}$,
$\langle({\nabla}{\mathbf{u}}_{i})\nu,({\nabla}{\mathbf{u}}_{j})\nu\rangle=\frac{2}{3}\cdot\frac{\lambda_{1}({\Omega}^{*})}{|{\Omega}|^{\frac{1}{3}}}\delta_{i,j}=\frac{1}{\sigma_{{\Omega}^{*}}^{2}}\delta_{i,j}.$
In this expression, $\delta_{i,j}$ is the usual Kronecker symbol and ${\sigma_{{\Omega}^{*}}^{2}}=\frac{3}{2}\cdot\frac{|{\Omega}|^{\frac{1}{3}}}{\lambda_{1}({\Omega}^{*})}$.
However, since each ${\mathbf{u}}_{i}$ is divergence free we have $\langle({\nabla}{\mathbf{u}}_{i})\nu,\nu\rangle=0$ (see Remark 5). In particular, the multiplicity of the first eigenspace $E_{1}$ is at most 2 (we recall that we are in dimension 3 here).
Now, let us argue by contradiction and assume that $\lambda_{1}({\Omega}^{*})$ has multiplicity 2. Let $({\mathbf{u}}_{1},{\mathbf{u}}_{2})$ be an orthonormal basis of $E_{1}$. From the previous computation, we deduce that, for any $x\in\partial{\Omega}^{*}$, $\{\sigma_{{\Omega}^{*}}({\nabla}{\mathbf{u}}_{1})\nu,\sigma_{{\Omega}^{*}}({\nabla}{\mathbf{u}}_{2})\nu\}$ is an orthonormal basis of the tangent plane $T_{x}(\partial{\Omega})$. Let us now argue using a scheme of proof inspired by [25, Proof of Theorem 2] to reach a contradiction.
@@
where $E_{1}^{\perp}$ is the $L^{2}$-orthogonal subspace to $E_{1}$ and
$\omega=-\int_{\partial{\Omega}^{*}}\|({\nabla}{\mathbf{u}}_{i})\nu\|^{2}=-\mathrm{Per}(\partial{\Omega}^{*})\cdot\frac{2}{3}\cdot\frac{\lambda_{1}({\Omega}^{*})}{|{\Omega}|^{\frac{1}{3}}}.$ (30)
@@
$\mathcal{E}:W^{1,2}_{0}({\Omega})^{3}\ni{\mathbf{v}}\mapsto\frac{1}{2}\int_{{\Omega}^{*}}\|{\nabla}{\mathbf{v}}\|^{2}-\frac{\lambda_{1}({\Omega}^{*})}{2}\int_{{\Omega}^{*}}\|{\mathbf{v}}\|^{2}$
@@
$X_{i}:=\{{\mathbf{v}}\in W^{1,2}_{0}({\Omega})^{2}\,,{\nabla}\cdot{\mathbf{v}}=0\,,{\mathbf{v}}=-({\nabla}{\mathbf{u}}_{i})\nu\text{ on }\partial{{\Omega}^{*}}\}\cap E_{1}^{\perp}.$
@@
$-\Delta{\mathbf{v}}_{i}-\lambda_{1}({\Omega}^{*}){\mathbf{v}}_{i}\in(\{{\mathbf{v}}\in W^{1,2}_{0}({\Omega})^{3}\,,{\nabla}\cdot{\mathbf{v}}=0\}\cap E_{1}^{\perp}).$
@@
Turning back to the proof of Theorem 4, observe that
@@
Here again the fact that ${\mathbf{v}}_{2}$ is tangential is instrumental. Thus $({\nabla}{\mathbf{v}}_{2}){\mathbf{v}}_{1}$ is orthogonal to both ${\mathbf{v}}_{1}$ and ${\mathbf{v}}_{2}$, so that there exists a function $f$ satisfying
$({\nabla}{\mathbf{v}}_{2}){\mathbf{v}}_{1}=f\nu.$
@@
$({\nabla}{\mathbf{v}}_{1}){\mathbf{v}}_{2}=g\nu$
for some function $g$. Finally, letting $[X,Y]$ be the Lie bracket of two vector fields, we deduce that
$[({\nabla}{\mathbf{v}}_{1}){\mathbf{v}}_{2},({\nabla}{\mathbf{v}}_{2}){\mathbf{v}}_{1}]=(f-g)\nu.$
However, $({\nabla}{\mathbf{v}}_{2}){\mathbf{v}}_{1}$ and $({\nabla}{\mathbf{v}}_{1}){\mathbf{v}}_{2}$ are tangential, whence their Lie bracket also is, so that
$[({\nabla}{\mathbf{v}}_{1}){\mathbf{v}}_{2},({\nabla}{\mathbf{v}}_{2}){\mathbf{v}}_{1}]=0.$
In particular, this implies that $\partial{\Omega}^{*}$ admits a set of two pointwise orthogonal, commuting vector fields. This implies that the Riemann tensor associated with $\partial{\Omega}^{*}$ endowed with the induced metric is zero (we refer for instance to [34, Theorem 3.1.7]). In particular, the Gauß  curvature of $\partial{\Omega}^{*}$ is zero everywhere, which is impossible.

### v2

@@ 3.3 The ball satisfies strong second-order conditions for Problem (5)
$u_{1}^{\prime}$ $=$ $\left.-\sin\theta\frac{\partial\psi_{\Phi}^{\prime}}{\partial r}-\cos\theta\frac{\partial\psi_{\Phi}^{\prime}}{\partial\theta}\right|_{r=1}=\sin\theta\,\frac{j_{1,1}}{\sqrt{\pi}}\,\varphi(\theta)$
@@
$u_{2}^{\prime}$ $=$ $\left.\cos\theta\frac{\partial\psi_{\Phi}^{\prime}}{\partial r}-\sin\theta\frac{\partial\psi_{\Phi}^{\prime}}{\partial\theta}\right|_{r=1}=-\cos\theta\,\frac{j_{1,1}}{\sqrt{\pi}}\,\varphi(\theta)$
From these two equations, we immediately obtain
$\frac{\partial\psi_{\Phi}^{\prime}}{\partial r}=-\frac{j_{1,1}}{\sqrt{\pi}}\,\varphi(\theta)\qquad\text{and}\qquad\frac{\partial\psi_{\Phi}^{\prime}}{\partial\theta}=0.$
Now since
$\frac{\partial\psi_{\Phi}^{\prime}}{\partial r}=\sum_{n=0}^{+\infty}j_{1,1}J_{n}^{\prime}(j_{1,1})(a_{n}\cos(n\theta)+b_{n}\sin(n\theta))+n(c_{n}\cos(n\theta)+d_{n}\sin(n\theta))$
and
$\frac{\partial\psi_{\Phi}^{\prime}}{\partial\theta}=\sum_{n=0}^{+\infty}nJ_{n}(j_{1,1})(b_{n}\cos(n\theta)-a_{n}\sin(n\theta))+n(d_{n}\cos(n\theta)-c_{n}\sin(n\theta))$
we obtain, by identification the four relations
$j_{1,1}J_{n}^{\prime}(j_{1,1})a_{n}+nc_{n}=-\frac{j_{1,1}}{\sqrt{\pi}}\alpha_{n}$ (27)
$j_{1,1}J_{n}^{\prime}(j_{1,1})b_{n}+nd_{n}=-\frac{j_{1,1}}{\sqrt{\pi}}\beta_{n}$ (28)
$J_{n}(j_{1,1})b_{n}+d_{n}=0$ (29)
$J_{n}(j_{1,1})a_{n}+c_{n}=0$ (30)
We obtain the desired expression by solving the two systems in $a_{n},c_{n}$ and $b_{n},d_{n}$ respectively. ∎
@@
Recall that we assume $\varphi\in\langle 1\rangle^{\perp}$ where $1$ is the constant function and $\langle 1\rangle^{\perp}$ is its $L^{2}$ orthogonal subspace.
@@
(we recall that we have supposed $\varphi$ orthogonal to $1,\cos\theta,\sin\theta$, so $a_{0}=a_{1}=b_{1}=0$.) Since the first positive zero of $J_{n}$ is greater than $j_{1,1}$ for all $n\geqslant 2$, one has $\gamma_{n}>0$ for every $n\geqslant 2$. Furthermore, using that $J_{n}(x)\sim\frac{1}{n!}\left(\frac{x}{2}\right)^{n}$ as $n\to+\infty$, we infer that
@@ 4 Necessary optimality conditions: proof of Theorem 4
As alluded to in the introduction, Theorem 3 follows from Theorem 4, so that it is clearer to begin with the proof of the latter (nevertheless, we give a proof of Theorem 3 in section 5 that only relies on the computations of the semi-differential analysed in the present section). This theorem relies on several results on the semi-differentiability of eigenvalues. To write things down in a precise way, we consider a domain ${\Omega}$ whose boundary is of class $\mathscr{C}^{2,\alpha}$ with $\alpha\in(0,1)$. The solution ${\mathbf{u}}$ thus belongs to $C^{2,\alpha}(\overline{{\Omega}})$ by standard elliptic regularity. Although $\lambda_{1}({\Omega})$ may not be simple, it has finite multiplicity, say $N$, and the map $\mathcal{F}$ is semi-shape differentiable at ${\Omega}$ [20, Theorem 2.1, Chapter 1] (see also [13, Theorem 2.1 and Remark 2.2] and [15]), in the sense that, for any vector field $\Phi\in\mathscr{X}$, the limit
$\langle\partial\mathcal{F}({\Omega}),\Phi\rangle:=\underset{t\searrow 0}{\lim}\frac{\mathcal{F}((\mathrm{Id}+t\Phi){\Omega})-\mathcal{F}({\Omega})}{t}$
@@
$\langle\partial\lambda_{1}({\Omega}),\Phi\rangle=-\min_{{\mathbf{u}}\in E_{1}\,,\int_{\Omega}\|{\mathbf{u}}\|^{2}=1}\int_{\partial{\Omega}}\|({\nabla}{\mathbf{u}})\nu\|^{2}\langle\Phi,\nu\rangle.$
@@
$\langle\partial\mathcal{F}({\Omega}),\Phi\rangle$ $=\frac{2}{3}\lambda_{1}({\Omega})|{\Omega}|^{-\frac{1}{3}}\int_{\partial{\Omega}}\langle\Phi,\nu\rangle-|{\Omega}|^{\frac{2}{3}}\min_{{\mathbf{u}}\in E_{1}\,,\int_{\Omega}\|{\mathbf{u}}\|^{2}=1}\int_{\partial{\Omega}}\|({\nabla}{\mathbf{u}})\nu\|^{2}\langle\Phi,\nu\rangle$
$=\min_{{\mathbf{u}}\in E_{1}\,,\int_{\Omega}\|{\mathbf{u}}\|^{2}=1}\int_{\partial{\Omega}}\left(\frac{2}{3}\lambda_{1}({\Omega})|{\Omega}|^{-\frac{1}{3}}-|{\Omega}|^{\frac{2}{3}}\cdot\|({\nabla}{\mathbf{u}})\nu\|^{2}\right)\langle\Phi,\nu\rangle$
$=\min_{(\alpha_{i})_{i=1,\dots,N},\ \sum_{i=1}^{N}{\alpha_{i}^{2}}=1}\sum_{i,j=1}^{N}\int_{\partial{\Omega}}\left(\frac{2}{3}\lambda_{1}({\Omega})|{\Omega}|^{-\frac{1}{3}}\alpha_{i}^{2}-|{\Omega}|^{\frac{2}{3}}\alpha_{i}\alpha_{j}\langle({\nabla}{\mathbf{u}}_{i})\nu,({\nabla}{\mathbf{u}}_{j})\nu\rangle\right)\langle\Phi,\nu\rangle.$
@@
$M_{\Phi}:=\frac{2}{3}\lambda_{1}({\Omega})|{\Omega}|^{-\frac{1}{3}}\int_{\partial{\Omega}}\langle\Phi,\nu\rangle\mathrm{I}_{3}-|{\Omega}|^{\frac{2}{3}}\left(\int_{\partial{\Omega}}\langle({\nabla}{\mathbf{u}}_{i})\nu,({\nabla}{\mathbf{u}}_{j})\nu\rangle\langle\Phi,\nu\rangle\right)_{1\leqslant i,j\leqslant N}.$
@@
We thus obtain the following fact: if ${\Omega}^{*}$ is optimal then for any $(i,j)\in\{1,\dots,N\}^{2}$, for any $x\in\partial{\Omega}^{*}\color[rgb]{0,0,0}$,
$\langle({\nabla}{\mathbf{u}}_{i})\nu,({\nabla}{\mathbf{u}}_{j})\nu\rangle=\frac{2}{3}\cdot\frac{\lambda_{1}({\Omega}^{*})}{|{\Omega}^{*}|}\delta_{i,j}=\frac{1}{\sigma_{{\Omega}^{*}}^{2}}\delta_{i,j}.$
In this expression, $\delta_{i,j}$ is the usual Kronecker symbol and ${\sigma_{{\Omega}^{*}}^{2}}=\frac{3}{2}\cdot\frac{|{\Omega}^{*}|}{\lambda_{1}({\Omega}^{*})}$.
However, since each ${\mathbf{u}}_{i}$ is divergence free we have $\langle({\nabla}{\mathbf{u}}_{i})\nu,\nu\rangle=0$ (see Remark 7). In particular, the multiplicity of the first eigenspace $E_{1}$ is at most 2 (we recall that we work in dimension 3).
Now, let us argue by contradiction and assume that $\lambda_{1}({\Omega}^{*})$ has multiplicity 2. Let $({\mathbf{u}}_{1},{\mathbf{u}}_{2})$ be an orthonormal basis of $E_{1}$. From the previous computation, we deduce that, for any $x\in\partial{\Omega}^{*}$, $\{\sigma_{{\Omega}^{*}}({\nabla}{\mathbf{u}}_{1})\nu,\sigma_{{\Omega}^{*}}({\nabla}{\mathbf{u}}_{2})\nu\}$ is an orthonormal basis of the tangent plane $T_{x}(\partial{\Omega})$. Let us now argue using a scheme of proof inspired by [28, Proof of Theorem 2] to reach a contradiction.
@@
where $E_{1}^{\perp}$ is the $L^{2}$-orthogonal subspace to $E_{1}$ and the constant $\omega$ is defined by
$\omega=-\int_{\partial{\Omega}^{*}}\|({\nabla}{\mathbf{u}}_{i})\nu\|^{2}=-\frac{2}{3}\mathrm{Per}(\partial{\Omega}^{*})\cdot\frac{\lambda_{1}({\Omega}^{*})}{|{\Omega}^{*}|}.$ (34)
@@
$\mathcal{E}:W^{1,2}_{0}({\Omega};\mathbb{R}^{3})\color[rgb]{0,0,0}\ni{\mathbf{v}}\mapsto\frac{1}{2}\int_{{\Omega}^{*}}\|{\nabla}{\mathbf{v}}\|^{2}-\frac{\lambda_{1}({\Omega}^{*})}{2}\int_{{\Omega}^{*}}\|{\mathbf{v}}\|^{2}$
@@
$X_{i}:=\{{\mathbf{v}}\in W^{1,2}_{0}({\Omega};\mathbb{R}^{3})\color[rgb]{0,0,0}\,,{\nabla}\cdot{\mathbf{v}}=0\,,{\mathbf{v}}=-({\nabla}{\mathbf{u}}_{i})\nu\text{ on }\partial{{\Omega}^{*}}\}\cap E_{1}^{\perp}.$
@@
$-\Delta{\mathbf{v}}_{i}-\lambda_{1}({\Omega}^{*}){\mathbf{v}}_{i}\in\left(\{{\mathbf{v}}\in W^{1,2}_{0}({\Omega};\mathbb{R}^{3})\color[rgb]{0,0,0}\,,{\nabla}\cdot{\mathbf{v}}=0\}\cap E_{1}^{\perp}\right).$
@@
Turning back to the proof of Theorem 4, let us write ${\mathbf{v}}_{i}=(v_{i,k})_{1\leqslant k\leqslant 3}$ for $i=1,2$ and consider the scalar function $\langle{\mathbf{v}}_{1},{\mathbf{v}}_{2}\rangle$ defined by $\langle{\mathbf{v}}_{1},{\mathbf{v}}_{2}\rangle(\cdot)=\sum_{k=1}^{3}v_{1,k}(\cdot)v_{2,k}(\cdot)$.
Then, denoting by ${\nabla}(\langle{\mathbf{v}}_{1},{\mathbf{v}}_{2}\rangle)$ the gradient of $\langle{\mathbf{v}}_{1},{\mathbf{v}}_{2}\rangle$ and still retaining, with a slight abuse of notations, the notation ${\nabla}{\mathbf{v}}_{i}$ for the Jacobian matrix of ${\mathbf{v}}_{i}$, observe that
@@
Here again the fact that ${\mathbf{v}}_{2}$ is tangential is instrumental. Thus $({\nabla}{\mathbf{v}}_{2}){\mathbf{v}}_{1}$ is orthogonal to both ${\mathbf{v}}_{1}$ and ${\mathbf{v}}_{2}$, so that there exists a function $f:\partial{\Omega}\to\mathbb{R}$ satisfying
${({\nabla}{\mathbf{v}}_{2}(x)){\mathbf{v}}_{1}(x)=f(x)\nu(x),\qquad x\in\partial{\Omega}.}$
@@
${({\nabla}{\mathbf{v}}_{1}(x)){\mathbf{v}}_{2}(x)=g(x)\nu(x),\qquad\text{a.e. }x\in\partial{\Omega}}$
for some function $g$. Finally, letting $[X,Y]$ be the Lie bracket of two vector fields,
${[{\mathbf{v}}_{1},{\mathbf{v}}_{2}]=({\nabla}{\mathbf{v}}_{1}){\mathbf{v}}_{2}-({\nabla}{\mathbf{v}}_{2}){\mathbf{v}}_{1}=(f-g)\nu\qquad\text{on }\partial{\Omega}.}$
However, as ${\mathbf{v}}_{1}$ and ${\mathbf{v}}_{2}$ are both tangential, so is their Lie bracket, whence $[{\mathbf{v}}_{1},{\mathbf{v}}_{2}]=0$. In particular, this implies that $\partial{\Omega}^{*}$ admits a set of two pointwise orthogonal, commuting vector fields. This implies that the Riemann tensor associated with $\partial{\Omega}^{*}$ endowed with the induced metric is zero (we refer for instance to [39, Theorem 3.1.7] or [42, Theorem 7.3 and its proof]). In particular, the Gauß  curvature of $\partial{\Omega}^{*}$ is zero everywhere, which is impossible.